QUESTIONI Titlow, Inc., produces and sells a single product. The product sells for $220.00 per unit and its variable expense is S57.20 per unit. The companys monthly fixed expense is $713,064. Required: 1. Calculate the Net operating Income of the company, if the 2. Determine the monthly break-even in units. Show your 3. If next month, the company expects to sell 4,000 units, do units sold are 5,000 work! you expect the company generating a profit or incurring a loss? Of how much? Explain. 4. Consider each of the following cases independently: a) If the management expects sales to increase by $40,000 because of an increase in the advertising budget by $18,000, how much will the profit increase by? Explain by showing your work. b) The sales manager suggests renting an equipment for $6,750 a month that will help in increasing the production, hence unit sales by 12%, but will also lead to an increase in each units selling price by 5%. Should the sales manager suggestion be accepted? Show your computations.
QUESTION II: Stanger Inc. produces and sells two products. Data concerning those products for the most recent month appear below: Product NI6S Product X07D $14,000 S6,720 $27,000 $12,550 Fixed expenses for the entire company were $17,570 Required: 1. Determine the overall break-even sales for the company. Show your work! 2. If the sales mix shifts toward Product N16S with no change in total sales, what will happen to the break-even sales for the company? Explain.
Answer #1

CM per unit = Selling price - Variable cost per unit.

CM per unit = $ 220 - $ 57.20 = $ 162.80.

Q-1.

Units sold = 5000.

CM = 5000 * 162.80 = $ 814000.

Profit = CM - Fixed Cost.

Profit = $ 814000 - $ 713064 = $ 100936.

Q-2

Break even units = ( Fixed cost ) / CM per unit

Break even units = $ 713064 / $ 162.80 = 4380 units.

Q-3.

If company sells 4000 units, company will incurr loss. Because company selling below break even point.

Profit / ( loss ) = 4000 * 162.80 - 713064 = $ 61864.

Extra sales units = 40000 / 220 =